# Learning Autonomous Surgical Irrigation and Suction with the da Vinci Research Kit Using Reinforcement Learning

### v1

@@ III-C Learning Environments for Irrigation and Suction
Based on CRESSim-ML, two simulated learning scenes are built for training agents to complete the irrigation and suction tasks autonomously and to achieve sim-to-real transfer. The overall configurations of both scenes are similar, consisting of a PSM robot with the Suction/Irrigator, a tissue container where fluids are present, a camera for generating image observations, and a background tabletop, as shown in Fig. 7a. To ensure the trained policy can be applied to the real-world setup, the simulated scenes are designed to resemble the real world, as will be discussed in Section V.
@@ III-C1 Irrigation Learning Environment
In each episode, a random tissue shape is generated in the container, as will be elaborated in Section IV-A. A random amount of fluid block with high cohesion and friction and a dark red color that simulates a clotted blood area is dropped onto the tissue. The goal is for the agent to control the PSM’s joint-space motion, move the end-effector (EE) toward the dense blood area, and irrigate liquid with lighter color and less dense fluid properties. When both types of liquid mix with each other, colors diffuse between particles as discussed in Section III-A. Additionally, each particle is initially assigned a value indicating whether it belongs to the clotted blood or the irrigation liquid. This value is also diffused across particles, providing information on the number of particles affected by irrigation. A threshold is used when determining whether the dense blood particles have been affected. After an enough number of blood particles are affected, the task is considered completed. A screenshot from the irrigation environment is shown in Fig. 7b.
@@
Observations and actions  The observation consists of both visual and vector parts. The visual part includes the RGB image from the camera and the initial frame of the camera, which provides information about the initial location of the dense blood. The vector observation includes current robot joint positions and a number indicating whether the robot is in contact with the tissue, which is decided by measuring the EE link’s incoming joint force and torque. The vector observation is stacked with values from the previous 3 steps to include historical motion data. The actions are the incremental movements of the robot joints, along with a control signal to toggle the irrigation on and off. The action frequency is 10 Hz.
@@
Reward function  The reward function is composed of several parts, including a reward proportional to the number of particles affected by irrigation at each step, a reward when the EE moves horizontally closer to the blood, and a completion reward. A reward is also given for activating irrigation when the EE is close to the blood, and deactivating it when they are not close. Conversely, it penalizes activating irrigation when distant or deactivating it when near. This encourages the agent to only turn on the irrigation while the EE is close to the blood. Without these, the agent can easily learn a suboptimal solution of always turning on irrigating, which is not a desired behavior in practice. Additionally, penalties are given when the EE’s orientation deviates from being vertical, and when the EE is in contact with the tissue, to encourage a safer behavior of the agent in the real world. For clarity, the reward function can be expressed by the summation of the reward features $\psi^{r}_{i}$ multiplied by their weights $w^{r}_{i}$:
@@
as listed in Table I.
@@ III-C2 Suction Learning Environment
The suction task considers removing the liquid from the container using the Suction/Irrigator, a step typically following irrigation. In the simulated learning environment, similar configurations to the irrigation environment are used. In each episode, a random number of fluid blocks with various amounts of liquid is dropped from different locations to the tissue container, forming a variety of initial liquid areas that simulate a mixture of blood, saline, and other possible contaminants. The goal is for the agent to control the PSM’s EE to suction all liquid in the container away. The suction model follows the implementation of [16], where a cone-shaped force field is applied around the EE of the PSM to move all particles in the region close to the force center. Particles that are close enough to the force center are removed from the scene. In this task, suction is always turned on and is not controlled by the agent for simplicity. A snapshot from the suction environment is shown in Fig. 7c.
Observations and actions  The observations and actions are very similar to those in the irrigation environment. The observation includes the RGB image from the camera and the stacked vector observation of current robot joint positions along with a number indicating whether the robot is in contact with the tissue. As the initial camera frame is, in principle, not necessary for making a decision, it is not provided as part of the observation in this task. The actions are the incremental movements of the robot joints without any additional control signals, as suction is always turned on.
Reward function  Same as in [16], the agent is rewarded when particles are removed from the tissue container. However, it is challenging to train the agent using this reward alone due to the sparsity of the reward, as it is likely that no liquid particles are removed most of the time. As more liquid is already removed, it is increasingly challenging for the agent to explore successful particle removal due to the few numbers left. This issue is signified by the existence of multiple different liquid areas caused by various tissue shapes and multiple random initial liquid blocks. To overcome this issue, the reward function also includes a reward when the EE is approaching the nearest liquid region. Task completion reward, EE orientation penalty, and contact penalty are also included, similar to the irrigation task. The reward components are listed in Table I.
IV Traning in the Simulator
@@ IV-A Environment Randomization and Curriculum Learning
Visual appearance  Visual aspects, including object colors, lighting, and camera pose are randomized in the simulation. After setting up the real-world configuration, as will be discussed in Section V, camera images are obtained from the real world. Based on the real-world images, the initial colors of various simulated objects are decided. During training, the object colors are randomly chosen from a range centering the initial values in the HSV color space. The objects with randomized colors include the Suction/Irrigator, the tissue in the container, the fluid, and the background tabletop. Similarly, an estimation of the camera pose in the real world is used to provide an initial value in the simulator, based on which randomness is added to the 3D position and 3 rotational angles during training. A directional light is used in the simulator, and its pose is randomized in a similar manner. The intensity and shadow strength of the light are also uniformly randomized within a specified range.
@@
Task variation  Moreover, we would like the agent to be able to handle a wide variety of task variations, such as with different tissue shapes and different blood locations. Following our previous work [16], the tissue shape is randomly generated using a Bézier surface by manipulating the control point locations. The robot’s initial joint positions are randomized according to a pre-defined range that roughly places the EE randomly above the container. For both irrigation and suction tasks, we randomize the initial amount of blood or liquid added to the scene. As discussed in Section III-C, the blood is split into different blocks dropped from different initial locations for the suction task.
@@
Physics  Similar to [16], we additionally randomize the physics properties, such as the cohesion and viscosity parameters of the irrigation liquid in the irrigation task and the liquid in the suction task. This provides diverse fluid behavior when irrigating and suctioning.
@@
Curriculum learning (CL) is a common technique to break down a complex task into simpler, more manageable stages and is a well-established method in RL [41]. We further utilize CL when training the irrigation agent by designing a two-lesson curriculum. In the first lesson, the irrigation reward is turned off and the agent only learns to approach the blood area without turning on the irrigation, and then turn it on when the EE is close enough to the blood. Additionally, we do not end the episode when enough particles are affected by irrigation, and no task completion reward is provided. In the second lesson, we restore the original task rewards to let the agent learn the actual task. Between these two lessons, we include transitional lessons where both task configurations (with and without the irrigation reward) are randomly sampled according to a proportion. Overall, the proportion of the actual task increases, gradually guiding the agent toward mastering the complete task through CL. Its effectiveness will be discussed in Section VI-A.
@@ IV-B Learning from Demonstration
where $\mathcal{D}_{E}$ is the expert demonstration dataset, $\mathbf{s}$ and $\mathbf{a}$ are the state and action, respectively. $\ell\left(\pi(\mathbf{s}),\,\mathbf{a}\right)$ is a discrepancy measure between the policy-predicted action $\pi(s)$ and the expert action $a$. A mean squared error (MSE) loss is used for continuous actions and a cross-entropy loss is used for discrete actions.
@@
Both BC and GAIL are commonly used with RL to take advantage of expert demonstrations and are implemented as part of ML-Agents. During the initial stage of training, regular RL optimization steps are followed by additional BC steps using $\mathcal{L}_{BC}$ with a separate optimizer. The learning rate decays linearly to zero, after which no BC steps will be carried out. The auxiliary reward from GAIL is added to the environment reward throughout training.
Figure 9: Representative photo of collecting human demonstrations.
To collect demonstrations with different characteristics and to reduce human effort, two types of demonstrations are collected through both scripted policies and human teleoperation in the simulated environment. When implementing the scripted policies, the actual particle states are used. For irrigation, all dense blood particles are looped over to find the horizontal center of the blood region, and the target position of the PSM’s EE is set to a point higher than it. Irrigation is turned off if they are not close enough, otherwise, it is turned on. For suction, all particles in the container are clustered into a few regions by building a spatial hash grid, similar to that discussed in Section III-A. The EE’s target position is set to a point higher than the medoid of the nearest blood cluster. For both scripted policies, the target orientation of the EE is always vertical. After obtaining the target EE pose, the Jacobian-based iterative inverse kinematics method is used to calculate the current-step joint-space action.
Similarly, when collecting human demonstrations, the pose of the real-world MTM robot’s EE is captured and set as the target EE pose of the PSM, from which the per-step joint-space actions are calculated using inverse kinematics. In addition, the MTM gripper pinch signal is used to indicate whether irrigation is on or off for the irrigation task. Fig. 9 shows how human demonstrations are collected.
@@ IV-C Training Configurations
The physics simulation (environment) step size is 0.02 seconds. However, the decision (action) frequency is lower than the simulation steps at once per 0.1 seconds, resulting in a decision frequency of 10 Hz. In the initial stage of each episode, the PSM robot is driven to the random initial position, and the fluids are dropped from a height into the container, during which no actions will be taken. The maximum number of allowed environment steps are 1,000 and 2,000 for irrigation and suction tasks, respectively. The episode terminates early if the task is considered complete, as described in Section III-C.
@@ V Experimental Setup
In the experiments, we first evaluate the performance of the irrigation and suction agents separately on their respective tasks. Then, we assess the overall performance of the complete two-step task, which involves irrigating first and then suctioning. The robot is controlled at less than 10 Hz in the real world.
@@ VI-A Training Results
For irrigation, we conduct training in four distinct configurations: with and without the curriculum, and with and without IL. This allows us to compare the impact of both CL and IL individually and in combination. In the case of suction, we only train the agent in two different settings: with and without IL. Fig. 12 shows the learning curves. It is noticed that for irrigation, the designed curriculum helps achieve better training performance. Conversely, IL does not help the training process for irrigation, in both cases with and without using CL. Instead, adding IL resulted in a counter effect on training. However, adding IL resulted in a better learning curve for suction. The negative effect of IL for irrigation suggests a misalignment between the GAIL reward and the actual task reward, possibly because the expert policies are not aligned closely with the optimal policy that can be learned in irrigation. This can be evidenced by comparing the agents’ performance with the human baselines. The human baseline is much higher than the achieved return by the irrigation agents, whereas they are relatively close in suction.
@@ VI-B Real-World Performance of Irrigation Only
We are not able to directly compare the real-world performance with the evaluation results in the simulator, as we cannot measure the true amount of ketchup being affected by irrigation. However, it appears that the real-world performance is slightly inferior to the one in the simulator in terms of accurately targeting the EE toward the ketchup before irrigation. This is likely due to the sim-to-real gap that will be discussed in Section VII-A. Snapshots from autonomous irrigation are shown in Fig. 15.
@@
Figure 14: Results from real-world trials for autonomous irrigation and suction.
Figure 15: Snapshots from real-world irrigation experiments.
Figure 16: Snapshots from real-world suction experiments.
Figure 17: Snapshots from complete irrigation-suction experiments in the real world.

### v2

@@ III-C Learning Environments for Irrigation and Suction
Based on CRESSim-ML, two simulated learning scenes are built for training agents to complete the irrigation and suction tasks autonomously and to achieve sim-to-real transfer. The overall configurations of both scenes are similar, consisting of a PSM robot with the Suction/Irrigator, a tissue container where fluids are present, a camera for generating image observations, and a background tabletop, as shown in Fig. 6a. To ensure the trained policy can be applied to the real-world setup, the simulated scenes are designed to resemble the real world, as will be discussed in Section V. It is worth noting that the environment setup is not intended to realistically represent a specific surgical procedure but rather to validate the automation of the surgical task considered in this work.
@@
In this work, we treat irrigation and suction as two sub-tasks completed by two separate agents, instead of training a unified agent that completes the full process. A key challenge in training a unified agent is defining an effective reward function. While the purpose of suction is well-understood, the goal of irrigation is usually ambiguous when described mathematically. If the reward is based solely on complete contaminant removal, the agent will learn to prioritize suction alone, even though irrigation is beneficial for making the contaminant less sticky. In fact, without irrigation, it is still possible to suction and remove most contaminants in practice, as long as the Suction/Irrigator is in close contact with the contaminant. However, irrigation not only aids in dilution but also prevents contaminants from sticking within the Suction/Irrigator. Therefore, considering two separate sub-tasks is a more natural choice. Moreover, separating the agents provides practical flexibility. For instance, in applications like blood suction (as explored in previous studies, e.g., [17, 18, 19]), our suction agent could be deployed independently without the irrigation phase.
It is also worth noting that in general, we consider models that observe a vision input from the camera and a vector input of the robot joint, and output the joint-space position increment. Since the action is in the joint space, the observation of the current joint positions is needed. This allows direct mapping to low-level controller commands in the joint space without requiring an inverse kinematic solver. In this work, using Cartesian end-effector (EE) position increments as agent actions is also impractical, as the PSM combined with the Suction/Irrigator has only 5 DoFs. A Cartesian space action requires 6-DoF position increments, but since the robot is kinematically constrained, it lacks the necessary DoFs to support this approach. For example, if the agent’s output specifies pure translation of the end-effector along the X-axis while keeping all other DoFs at zero, the system will typically introduce an unintended rotational component. This occurs because the robot has only 5 DoFs, making it kinematically impossible to achieve arbitrary 6-DoF Cartesian motion. As a result, the robot moves in a way that satisfies its own constraints, violating the agent’s intended zero-motion in other DoFs. Ideally, there should be no significant performance difference between joint-space and Cartesian actions, provided the current robot position is included in the observation. Furthermore, we employ incremental actions rather than absolute joint positions, as is common in RL-based control work, such as in [4, 29, 3, 36]. Incremental actions mitigate training challenges caused by absolute-value scaling, where small workspace variations become difficult to learn. Additionally, they help maintain consistent execution timing, preventing large, stepwise motions that could disrupt inference-time control consistency.
@@ III-C1 Irrigation Learning Environment
In each episode, a random tissue shape is generated in the container, as will be elaborated in Section IV-A. A random amount of fluid block with high cohesion and friction and a dark red color that simulates a clotted blood area is dropped onto the tissue. The goal is for the agent to control the PSM’s joint-space motion, move the EE toward the dense blood area, and irrigate with liquid of lighter color and less dense fluid properties. When both types of liquid mix with each other, colors diffuse between particles as discussed in Section III-A. Additionally, each particle is initially assigned a value indicating whether it belongs to the clotted blood or the irrigation liquid. This value is also diffused across particles, providing information on the number of particles affected by irrigation. A threshold is used when determining whether the dense blood particles have been affected. Particles that spill outside of the tissue container will be removed and not considered, as they are set as inactive. After enough blood particles are affected, the task is considered completed. A screenshot from the irrigation environment is shown in Fig. 6b.
@@
Observations and actions  The observation consists of both visual and vector parts. The visual component includes an RGB image from the camera at each step, as well as an initial frame captured at the beginning of the episode. The initial frame provides implicit information about the dense blood’s starting location. This is crucial because, as more liquid is added, it becomes difficult to determine where the dense blood was initially, as the liquid spreads and covers the area. The vector observation includes current robot joint positions and a number indicating whether the robot is in contact with the tissue, which is decided by measuring the EE link’s incoming joint force and torque. The vector observation is stacked with values from the previous 3 steps to include historical motion data. The actions are the incremental movements of the robot joints, along with a control signal to toggle the irrigation on and off. The action frequency is 10 Hz.
@@
Reward function  The reward function is composed of several parts, including a reward proportional to the number of particles affected by irrigation at each step, a reward when the EE moves horizontally closer to the blood, and a completion reward. A reward is also given for activating irrigation when the EE is close to the blood, and deactivating it when they are not close. Conversely, it penalizes activating irrigation when distant or deactivating it when near. This encourages the agent to only turn on the irrigation while the EE is close to the blood. Without these, the agent can easily learn a suboptimal solution of always turning on irrigation, which is not a desired behavior in practice. Additionally, penalties are given when the EE’s orientation deviates from being vertical, and when the EE is in contact with the tissue, to encourage safer behavior of the agent in the real world. While the penalty related to EE’s orientation is always assigned, the joint limits of the robot are implemented and respected according to the real-world robot configuration, and may prevent the EE from achieving a fully vertical orientation in certain configurations. For clarity, the reward function can be expressed by the summation of the reward features $\psi^{r}_{i}$ multiplied by their weights $w^{r}_{i}$:
@@
as listed in Table I. These weighting values are selected based on the scale of the features, and training trials are conducted to tune them for ideal performance.
@@ III-C2 Suction Learning Environment
The suction task considers removing the liquid from the container using the Suction/Irrigator, a step typically following irrigation. In the simulated learning environment, similar configurations to the irrigation environment are used. In each episode, a random number of fluid blocks with various amounts of liquid are dropped from different locations into the tissue container, forming a variety of initial liquid areas that simulate a mixture of blood, saline, and other possible contaminants. The goal is for the agent to control the PSM’s EE to suction all liquid in the container away. The suction model follows the implementation of [19], where a cone-shaped force field is applied around the EE of the PSM to move all particles in the region close to the force center. While it may be appropriate to use the terminal states of the irrigation task as the starting point for the suction task, this approach could limit the range of scenarios the suction agent can handle. As mentioned earlier, our goal is to develop a general suction agent that is less dependent on the irrigation process. Particles that are close enough to the force center are removed from the scene. In this task, suction is always turned on and is not controlled by the agent for simplicity. A snapshot from the suction environment is shown in Fig. 6c.
Observations and actions  The observations and actions are very similar to those in the irrigation environment. The observation includes the RGB image from the camera and the stacked vector observation of current robot joint positions, along with a number indicating whether the robot is in contact with the tissue. As the initial camera frame is, in principle, not necessary for making a decision, it is not provided as part of the observation in this task. The actions are the incremental movements of the robot joints without any additional control signals, as suction is always turned on.
Reward function  Same as in [19], the agent is rewarded when particles are removed from the tissue container. However, it is challenging to train the agent using this reward alone due to the sparsity of the reward, as it is likely that no liquid particles are removed most of the time. As more liquid is already removed, it is increasingly challenging for the agent to explore successful particle removal due to the few number left. This issue is signified by the existence of multiple different liquid areas caused by various tissue shapes and multiple random initial liquid blocks. To overcome this issue, the reward function also includes a reward when the EE is approaching the nearest liquid region. Task completion reward, EE orientation penalty, and contact penalty are also included, similar to the irrigation task. The reward components are listed in Table I.
IV Training in the Simulator
@@ IV-A Environment Randomization and Curriculum Learning
Visual appearance  Visual aspects, including object colors, lighting, and camera pose, are randomized in the simulation. After setting up the real-world configuration, as will be discussed in Section V, camera images are obtained from the real world. Based on the real-world images, the initial colors of various simulated objects are decided. During training, the object colors are randomly chosen from a range centered on the initial values in the HSV color space. The objects with randomized colors include the Suction/Irrigator, the tissue in the container, the fluid, and the background tabletop. Similarly, an estimation of the camera pose in the real world is used to provide an initial value in the simulator, based on which randomness is added to the 3D position and 3 rotational angles during training. A directional light is used in the simulator, and its pose is randomized in a similar manner. The intensity and shadow strength of the light are also uniformly randomized within a specified range.
@@
Task variation  Moreover, we would like the agent to be able to handle a wide variety of task variations, such as with different tissue shapes and different blood locations. Following previous work [19], the tissue shape is randomly generated using a Bézier surface by manipulating the control point locations. The robot’s initial joint positions are randomized according to a pre-defined range that roughly places the EE randomly above the container. For both irrigation and suction tasks, we randomize the initial amount of blood or liquid added to the scene. As discussed in Section III-C, the blood is split into different blocks dropped from different initial locations for the suction task.
@@
Physics  Similar to [19], we additionally randomize the physics properties, such as the cohesion and viscosity parameters of the irrigation liquid in the irrigation task and the liquid in the suction task. This provides diverse fluid behavior when irrigating and suctioning, and helps bridge the gap between simulation and real-world dynamics.
@@
Curriculum learning (CL) is a common technique to break down a complex task into simpler, more manageable stages and is a well-established method in RL [47]. We further utilize CL when training the irrigation agent by designing a two-lesson curriculum. In the first lesson, the irrigation reward is turned off, and the agent only learns to approach the blood area without turning on the irrigation, and then turns it on when the EE is close enough to the blood. Additionally, we do not end the episode when enough particles are affected by irrigation, and no task completion reward is provided. In the second lesson, we restore the original task rewards to let the agent learn the actual task. Between these two lessons, we include transitional lessons where both task configurations (with and without the irrigation reward) are randomly sampled according to a proportion. Overall, the proportion of the actual task increases, gradually guiding the agent toward mastering the complete task through CL. Its effectiveness will be discussed in Section VI-A.
@@ IV-B Learning from Demonstration
where $\mathcal{D}_{E}$ is the expert demonstration dataset, $\mathbf{s}$ and $\mathbf{a}$ are the state and action, respectively. $\ell\left(\pi(\mathbf{s}),\,\mathbf{a}\right)$ is a discrepancy measure between the policy-predicted action $\pi(s)$ and the expert action $a$. A mean squared error (MSE) loss is used for continuous actions, and a cross-entropy loss is used for discrete actions.
@@
Both BC and GAIL are commonly used with RL to take advantage of expert demonstrations and are implemented as part of ML-Agents. During the initial stage of training, regular RL optimization steps are followed by additional BC steps using $\mathcal{L}_{BC}$ with a separate optimizer. The learning rate decays linearly to zero, after which no BC steps will be carried out. The auxiliary reward from GAIL is added to the environment reward throughout training. A similar approach has been previously investigated and shown success on tissue manipulation in [48].
To collect demonstrations with different characteristics and to reduce human effort, two types of demonstrations are collected through both scripted policies and human teleoperation in the simulated environment. When implementing the scripted policies, the actual particle states are used, although they are not observed by the agent. For irrigation, all dense blood particles are looped over to find the horizontal center of the blood region, and the target position of the PSM’s EE is set to a point higher than that. Irrigation is turned off if they are not close enough, otherwise, it is turned on. For suction, all particles in the container are clustered into a few regions by building a spatial hash grid, similar to that discussed in Section III-A. The EE’s target position is set to a point higher than the medoid of the nearest blood cluster. For both scripted policies, the target orientation of the EE is always vertical. After obtaining the target EE pose, the Jacobian-based iterative inverse kinematics method is used to calculate the current-step joint-space action.
Similarly, when collecting human demonstrations, the pose of the real-world MTM robot’s EE is captured and set as the target EE pose of the PSM, from which the per-step joint-space actions are calculated using inverse kinematics. In addition, the MTM gripper pinch signal is used to indicate whether irrigation is on or off for the irrigation task.
@@ IV-C Training Configurations
The physics simulation (environment) step size is 0.02 seconds. However, the decision (action) frequency is lower than the simulation steps at once per 0.1 seconds, resulting in a decision frequency of 10 Hz. In the initial stage of each episode, the PSM robot is driven to a random initial position, and the fluids are dropped from a height into the container, during which no actions will be taken. The maximum number of allowed environment steps is 1,000 and 2,000 for irrigation and suction tasks, respectively. The episode terminates early if the task is considered complete, as described in Section III-C.
@@ V Experimental Setup
In the experiments, we first evaluate the performance of the irrigation and suction agents separately on their respective tasks. Then, we assess the overall performance of the complete two-step task, which involves irrigating first and then suctioning. The robot is controlled at around 5 to 10 Hz in the real world, varying due to robot motion execution time at each step. We intentionally keep the robot motion slower than in training for safety. The obtained model is converted into the ONNX format with default optimizations by ML-Agents and inferred on Intel(R) Core(TM) i7-9700K. Image frames are captured from the webcam at each step in an asynchronous manner and fed into the model with the current robot joint state. The output actions are set as an incremental joint position target and sent to the dVRK software through its ROS package.
@@ VI-A Training Results
For irrigation, we conduct training in four distinct configurations: with and without the curriculum, and with and without IL. This allows us to compare the impact of both CL and IL individually and in combination. In the case of suction, we only train the agent in two different settings: with and without IL. Fig. 10 shows the learning curves. It is noticed that for irrigation, the designed curriculum helps achieve better training performance. Conversely, IL does not help the training process for irrigation, in both cases with and without using CL. Instead, adding IL resulted in a counter-effect on training. However, adding IL resulted in a better learning curve for suction. The negative effect of IL for irrigation suggests that the reward inferred by GAIL does not fully align with the main RL objective. This misalignment may arise because the learned GAIL reward signal may not fully guide toward the task-specific objective, or because the agent model cannot capture all of the required features, potentially leading to different optimization directions.
@@ VI-B Real-World Performance of Irrigation Only
We are not able to directly compare the real-world performance with the evaluation results in the simulator, as we cannot measure the true amount of ketchup being affected by irrigation. However, it appears that the real-world performance is slightly inferior to the one in the simulator in terms of accurately targeting the EE toward the ketchup before irrigation. Snapshots from autonomous irrigation are shown in Fig. 13.
@@
Figure 12: (a) Experimental procedure and (b) results for real-world combined irrigation-suction. Lines of different colors in (b) represent individual trials.
Figure 13: Representative snapshots selected at different action steps from real-world irrigation experiments.
Figure 14: Representative snapshots selected at different action steps from real-world suction experiments.
Figure 15: Representative snapshots selected at different action steps from complete irrigation-suction experiments in the real world.